# Exploring Local Memorization in Diffusion Models via Bright Ending Attention

### v1

@@ 5.1 Incorporating the bright ending (BE)
BE’s success in the new task of extracting localized memorized regions can also benefit existing tasks of evaluating, detecting, and mitigating memorization in diffusion models. By integrating the extracted local mask into current strategies, these approaches can adopt a local perspective, thereby reducing the performance gap associated with local memorization.
@@
$d=\frac{1}{T}\sum_{t=1}^{T}\left\|\left(\varepsilon_{\theta}(x_{t},e_{p})-\varepsilon_{\theta}(x_{t},e_{\phi})\right)\circ\mathbf{m}\right\|_{2}\bigg{/}\left(\frac{1}{N}\sum_{i=1}^{N}m_{i}\right)$ (8)
@@
$LS(\hat{x},x)=-\mathbbm{1}_{\text{SSCD}>0.5}\cdot\left\|(\hat{x}-x)\circ\mathbf{m}\right\|_{2}$ (9)
$S(\hat{x},x)=-\mathbbm{1}_{\text{SSCD}>0.5}\cdot\left\|(\hat{x}-x)\right\|_{2}$ (10)
@@ 5.2 Experiments
We follow the baseline methodology from Wen et al. (2024) and evaluate performance using the area under the curve (AUC) of the receiver operating characteristic (ROC) curve, the True Positive Rate at 1% False Positive Rate (T@1%F), and the F1 score, and report the detection performance during the 1, 10, and 50 inference steps. However, the baseline detection performance measures whether a prompt is prone to memorization rather than if a specific generation is memorized. They experimented with how well the average magnitude of 1, 4, and 32 generations per prompt differentiates between memorized and non-memorized cases. This approach is flawed because using different random seeds, a prompt prone to memorization can sometimes generate non-memorized images. Consequently, a prompt that rarely yields memorization might still be classified as memorized. For example, generating four non-memorized images from a prompt classified as memorized due to its occasional memorized output can falsely inflate performance metrics. Conversely, failing to classify these as non-memorized when they are all non-memorized should not be penalized. To address this, we assess detection performance based on accurately predicting each generation as memorized or not rather than each prompt. We re-implement the baseline method under the same conditions for a fair comparison. For mitigation, the output utility is evaluated based on the CLIP score, which measures the text-alignment of generated outputs. We evaluate similarity scores using both traditional SSCD and our proposed localized metric to provide a comprehensive performance view, experimenting with multiple privacy levels. Due to current metrics’ failure to accurately identify memorized cases, we manually label the ground truth for each generated image from the memorized prompts. Webster (2023) organizes the nearest training images for these generations, making it easy to observe local memorization cases with minimal subjectivity in manual labeling. We compare different criteria’s F1 scores, using a threshold of 0.50 for SSCD to determine memorized cases and a threshold of -50 for LS and S metrics as defined in Eq. 10 and Eq. 9. We separately report performance for local and global memorization in Tab. 1 and Fig. 7 and 7, highlighting how our localized strategies, after incorporating BE, effectively reduce the performance gap specifically caused by local memorization.
@@
Table 1: Detection performance for local and global memorization at different inference steps.
First Step First 10 Steps All Steps
@@
Baseline - Local 0.918 0.864 0.629 0.989 0.982 0.953 0.990 0.983 0.560
Ours - Local 0.943 0.893 0.731 0.995 0.987 0.985 0.996 0.988 0.926
Baseline - Global 0.979 0.944 0.934 1.000 0.987 1.000 0.999 0.976 1.000
Ours - Global 0.981 0.948 0.940 1.000 0.987 1.000 0.999 0.977 1.000
@@ Results.
Tab. 1 demonstrates that incorporating the memorization mask into magnitude computation effectively enhances detection performance across all metrics and various inference steps for local memorization while resulting in almost identical results for global memorization. Additionally, Fig. 3 and 5 (Right) further demonstrate this targeted improvement in local memorization. The magnitudes for non-memorization and global memorization remain almost unchanged, but the magnitudes for local memorization are successfully amplified, making them more distinguishable from non-memorization cases. This effectively reduces the performance gap caused by local memorization by converting previous False Negative cases into True Positives (examples shown in Fig. 8).
@@
For mitigation, we also include the original Stable Diffusion (SD) without any mitigation strategies applied in Fig. 7 and 7 for comparison, illustrating the extent to which we have reduced the local memorization performance gap. Specifically, for local memorization, under the high-utility scenario where we aim to minimally reduce the CLIP score while mitigating memorization, our approach achieves a significantly better trade-off than the baseline. We reduce SD’s similarity score by twice as much as the baseline while simultaneously observing a much smaller reduction in the CLIP score - less than half compared to the baseline. We also tested different privacy levels and consistently achieved a better trade-off for local memorization. Specifically, we observe a lower similarity score for the same CLIP score, as measured by both SSCD and our localized metric. Additionally, we consistently observe a higher CLIP score for the same similarity score. As expected, the results for global memorization remain almost identical to the baseline.
@@
It is worth noting that the effectiveness of these strategies not only demonstrates our improvement over the current state-of-the-art in existing tasks after incorporating BE’s local memorization mask, but also serves as an extrinsic validation of BE’s effectiveness in the new task we propose - extracting local memorization masks. To our knowledge, BE is the first method to precisely extract local memorization masks in a highly efficient manner, requiring only one inference step without the need to query the training data.
@@
Figure 8: Examples of converting FN into TP. Generated image followed by training image.
@@ 6 Related work
Webster (2023) categorizes memorized captions into three types: (1) Matching Verbatim (MV), where the generated image using a memorized caption is an exact replica of the corresponding image; (2) Retrieval Verbatim (RV), where the generated image using a memorized caption replicates other training images rather than the corresponding one; and (3) Template Verbatim (TV), where the generated image using a memorized caption replicates only the template shared by a subset of training images, with variations in fixed locations outside the template, corresponding to different captions. This categorization focuses on the causes of memorization: MV results from overfitting in the one-to-one mapping between image and caption in the training set, while many-to-many mappings lead to RV and TV. However, the concepts of Local Memorization (LM) and Global Memorization (GM) differ, focusing on whether the generated image exactly replicates an entire training image or just a portion of it, regardless of whether the training image matches the prompt’s corresponding image or the causes of memorization. RV and TV can be seen as subsets of LM, as there are additional causes of LM beyond the “many-to-many mapping overfitting” that remain unidentified. For instance, in Fig. 8, the image in the third column is classified as MV by Webster (2023) due to the one-to-one mapping of the training image-caption, rather than many-to-many. However, it only memorizes a local portion of the image in the last column. This suggests that a pre-trained model may occasionally exhibit creativity when recalling overfitted memories, which can also lead to local memorization.
@@
Carlini et al. (2023) detects memorization by calculating the Euclidean distance between generated images and training images, recommending re-training on de-duplicated data as a mitigation strategy.Daras et al. (2023) suggests mitigating memorization by training models on corrupted data. Somepalli et al. (2023b) proposes to mitigate memorization by injecting randomness into text prompts, such as adding random tokens or noise, to break the link between memorized prompts and their corresponding images. Chen et al. (2024) employs a nearest neighbor search, similar to Carlini et al. (2023), to identify potential memorized outputs and then selectively applies guidance methods to instruct these generations away from the memorized training images, demonstrating outstanding mitigation performance. Recently, Wen et al. (2024) proposed an efficient method for detecting memorization during the inference phase of diffusion models. This approach uses the magnitude of text-conditional predictions for prompt detection, enabling high accuracy from the initial inference step, and triggers a mitigation strategy to optimize the prompt embedding when memorization is detected. Both the proposed detection and mitigation strategies have achieved state-of-the-art results. A most recent work (Ren et al., 2024) utilized cross-attention patterns to design detection and mitigation strategies, achieving performance comparable to Wen et al. (2024). The key insight is that the cross-attention scores for memorized prompts are less concentrated across all tokens in a prompt during multiple inference steps. This lack of concentration is quantified using entropy, which captures the distribution of cross-attention at the prompt-level. While both leverage cross-attention, it is important to note that our bright ending cross-attention (BE) differs from the cross-attention entropy approach regarding objectives, insights, inputs, and the final computed statistic. Specifically, different from Ren et al. (2024)’s objective of leveraging cross-attention for designing detection and mitigation strategies, BE also aims to achieve a new task of precisely and efficiently locating localized memorization regions. Also, BE does not rely on the concentration patterns of all tokens during multiple steps as the insight and inputs but instead uses the raw final step attention score of the end token. Additionally, BE employs the raw image-patch level distribution rather than the prompt-level distribution as the computed statistic, serving as a local mask to highlight areas of local memorization.
7 Conclusion and limitations
Conclusion.
@@
In this paper, we identified a performance gap in existing tasks of measuring, detecting, and mitigating memorization in diffusion models, particularly in cases where the generated images memorize only parts of the training images. To address this, we introduced a new localized view of memorization in diffusion models, distinguishing between global and local memorization, with the latter being a more meaningful and generalized concept. To facilitate investigation from a local perspective, we pioneered a new task: extracting localized memorized regions. We proposed leveraging the newly observed ‘bright ending’ (BE) phenomenon to precisely and efficiently accomplish this task, resulting in a local memorization mask. Through extensive experiments, we demonstrated that the mask generated by this new task can also be used to enhance performance in existing tasks. Specifically, we proposed a simple yet effective strategy to incorporate the BE local memorization mask into existing frameworks, refining them to adopt a local perspective. This approach reduces the performance gap associated with local memorization and achieves new state-of-the-art results. The improved results not only highlight our improvements in existing tasks but also serve as extrinsic validation of BE’s effectiveness in the new task of local memorization mask extraction, underscoring the significant contribution of BE.
Limitations and future work.
One limitation of using the memorization mask extracted through the bright ending phenomenon in tasks like detection and mitigation is that it can only be observed at the end of the inference process. This requirement results in a longer processing time than existing strategies when applied to existing tasks such as detection and mitigation. However, we believe the additional time is justified by its significant positive contributions to investigating local memorization in diffusion models. Moreover, each inference takes only a few seconds on a consumer-grade GPU, making the process practical for real-world applications. While existing tasks are well-studied with established evaluation criteria, future work could explore how to evaluate the performance of the newly proposed local memorization region extraction task. Our current approach uses an extrinsic evaluation method that demonstrates improvements by incorporating the mask into existing strategies. Future research could also consider an intrinsic evaluation by directly comparing the extracted mask with ground truth masks. Another future direction could involve developing more advanced evaluation metrics for local memorization. In our work, we employed a straightforward approach by combining a masked version of the L2 distance with the widely adopted SSCD to compute similarity. Future research could focus on creating methods similar to SSCD, trained in a self-supervised manner but specifically tailored for more accurate localized copy detection.

### v2

@@ 5.1 Incorporating the bright ending (BE)
BE’s success in extracting localized memorized regions can also benefit existing tasks of evaluating, detecting, and mitigating memorization in diffusion models. By integrating the extracted local mask into current strategies, these approaches can adopt a local perspective, thereby reducing the performance gap associated with local memorization. Specifically, as discussed in Sec. 4, the bright ending (BE) attention map inherently serves as an automatically extracted memorization mask. This mask directly utilizes the BE attention scores as its values, eliminating the need for thresholding to convert it into a binary mask of only 0s and 1s. By doing so, it avoids the added variability introduced by thresholding hyperparameters, which may be less effective under certain threshold settings, thereby improving robustness. The approach remains effective regardless of whether the memorized regions are subtle or abstract, as it solely relies on the pre-trained model’s overfitted memory during memorization to extract the mask. This eliminates the dependence on factors such as the degree of subtlety or abstraction of memorized regions or thresholding based on size or intensity of memorization. For a qualitative evaluation of the mask, please refer to Fig. 1 and Fig. 4, where brighter patches indicate higher attention scores and darker patches indicate lower ones.
@@
$LD=\frac{1}{T}\sum_{t=1}^{T}\left\|\left(\varepsilon_{\theta}(x_{t},e_{p})-\varepsilon_{\theta}(x_{t},e_{\phi})\right)\circ\mathbf{m}\right\|_{2}\bigg{/}\left(\frac{1}{N}\sum_{i=1}^{N}m_{i}\right)$ (8)
@@
$LS(\hat{x},x)=-\mathbbm{1}_{\text{SSCD}<0.5}\cdot\left\|(\hat{x}-x)\circ\mathbf{m}\right\|_{2},$ (9)
$S(\hat{x},x)=-\mathbbm{1}_{\text{SSCD}<0.5}\cdot\left\|(\hat{x}-x)\right\|_{2}$ (10)
@@ 5.2 Experiments
We follow the baseline methodology from Wen et al. (2024) and evaluate performance using the area under the curve (AUC) of the receiver operating characteristic (ROC) curve, the True Positive Rate at 1% False Positive Rate (T@1%F), and the F1 score, and report the detection performance during the 1, 10, and 50 inference steps. However, the baseline detection performance measures whether a prompt is prone to memorization rather than if a specific generation is memorized. They experimented with how well the average magnitude of 1, 4, and 32 generations per prompt differentiates between memorized and non-memorized cases. This approach is flawed because using different random seeds, a prompt prone to memorization can sometimes generate non-memorized images. Consequently, a prompt that rarely yields memorization might still be classified as memorized. For example, generating four non-memorized images from a prompt classified as memorized due to its occasional memorized output can falsely inflate performance metrics. Conversely, failing to classify these as non-memorized when they are all non-memorized should not be penalized. To address this, we assess detection performance based on accurately predicting each generation as memorized or not rather than each prompt. We re-implement the baseline method under the same conditions for a fair comparison. For mitigation, the output utility is evaluated based on the CLIP score, which measures the text-alignment of generated outputs. We evaluate similarity scores using both traditional SSCD and our proposed localized metric to provide a comprehensive performance view, experimenting with multiple privacy levels. Due to current metrics’ failure to accurately identify memorized cases, we manually label the ground truth for each generated image from the memorized prompts. Webster (2023) organizes the nearest training images for these generations, making it easy to observe local memorization cases with minimal subjectivity in manual labeling. Please check out Sec. 9.3 for a detailed illustration of the labeling process.
We separately evaluate performance for local and global memorization, contrasting our localized strategies with the baseline’s globalized strategies. Table 1 presents a comparison between our localized detection strategy, which uses LD as the detection signal (Eq. 8), and the baseline globalized detection strategy that relies on D (Eq. 7). Figures 7 and 7 highlight comparisons of localized mitigation, utilizing LD (Eq. 8) as both the triggering signal and loss function, with globalized mitigation strategies that use D (Eq. 7). Finally, Table 2 compares the localized similarity metric LS (Eq. 9) against S (Eq. 10) and SSCD in terms of F1-score, using memorization thresholds. For SSCD, we follow the standard threshold of 0.50, while for LS and S, we adopt a threshold of -50.
@@
Table 1: Detection performance for local and global memorization at different inference steps $T$.
$T=1$ $T=10$ $T=50$
@@
Baseline (D) - Local 0.918 0.864 0.629 0.989 0.982 0.953 0.990 0.983 0.560
Ours (LD) - Local 0.943 0.893 0.731 0.995 0.987 0.985 0.996 0.988 0.926
Baseline (D) - Global 0.979 0.944 0.934 1.000 0.987 1.000 0.999 0.976 1.000
Ours (LD) - Global 0.981 0.948 0.940 1.000 0.987 1.000 0.999 0.977 1.000
@@ Results.
As the baseline detection strategy in Eq. 7 incorporates a hyperparameter $T$, where the detection signal is defined as the average magnitude over $T$ inference steps, Tab. 1 evaluates performance across different values of $T$. This evaluation compares the baseline’s magnitude with our proposed masked/localized magnitude, incorporating the BE mask as outlined in Eq. 8 under such different choices of baseline’s hyperparameter values for demonstrating BE’s ablation effects.
@@
The results in Tab. 1 consistently demonstrate that integrating the BE mask improves detection performance under varying baseline hyperparameter settings across all metrics for local memorization while maintaining the performance for global memorization. These findings validate (1) the BE mask’s effectiveness in accurately identifying memorized regions and (2) the importance of adopting a localized perspective for studying memorization, which remains under-explored. Additionally, Fig. 3 and 5 (Right) further demonstrate this targeted improvement in local memorization. The magnitudes for non-memorization and global memorization remain almost unchanged, but the magnitudes for local memorization are successfully amplified, making them more distinguishable from non-memorization cases. This effectively reduces the performance gap caused by local memorization by converting previous False Negative cases into True Positives (examples shown in Fig. 8).
@@
For mitigation, we also include the original Stable Diffusion (SD) without any mitigation strategies applied in Fig. 7 and 7 for comparison, illustrating the extent to which we have reduced the local memorization performance gap. The horizontal axis (CLIP score) represents utility, while the vertical axis (similarity score) reflects privacy risks. More favorable privacy-utility trade-offs are indicated by points further toward the bottom right. This placement signifies lower privacy risks for the same utility level (same x-axis value) and higher utility for the same level of privacy risks (same y-axis value). Specifically, for local memorization, under the high-utility scenario where we aim to minimally reduce the CLIP score while mitigating memorization, our approach achieves a significantly better trade-off than the baseline. We reduce SD’s similarity score by twice as much as the baseline while simultaneously observing a much smaller reduction in the CLIP score - less than half compared to the baseline. We also tested different privacy levels and consistently achieved a better trade-off for local memorization. Specifically, we observe a lower similarity score for the same CLIP score, as measured by both SSCD and our localized metric. Additionally, we consistently observe a higher CLIP score for the same similarity score. As expected, the results for global memorization remain almost identical to the baseline. This further confirms BE’s effectiveness in addressing the performance gap introduced by local memorization in the task of mitigating memorization.
For evaluation metric, Tab. 2 demonstrates the superiority of our localized metric, LS (Eq. 9), which integrates SSCD, L2, and our BE mask, over the globalized metric, S (Eq. 10), which uses only SSCD and L2. This highlights the contribution of incorporating the BE mask. Furthermore, the results underscore the limitations of solely relying on SSCD, which performs well for global memorization but struggles with local cases. Notably, our design of LS, which incorporates SSCD, retains strong performance for global memorization while addressing the performance gap in local memorization, further validating the benefit of such a design.
@@
It is worth noting that the effectiveness of these strategies not only demonstrates our improvement over the current state-of-the-art in existing tasks after incorporating BE’s local memorization mask, but also serves as an extrinsic validation of BE’s effectiveness in extracting local memorization masks. To our knowledge, BE is the first method to precisely extract local memorization masks in a highly efficient manner, requiring only one inference pass without the need to query the training data.
@@
Figure 8: Examples of turning FN to TP via mitigation strategy. Generated image followed by training image.
Table 2: F1-score comparisons for different evaluation metrics.
Local Global
SSCD 0.940 1.000
S 0.991 1.000
LS (Ours) 0.995 1.000
@@ 6 Related work
Webster (2023) categorizes memorized captions into three types: (1) Matching Verbatim (MV), where the generated image using a memorized caption is an exact replica of the corresponding image; (2) Retrieval Verbatim (RV), where the generated image using a memorized caption replicates other training images rather than the corresponding one; and (3) Template Verbatim (TV), where the generated image using a memorized caption replicates only the template shared by a subset of training images, with variations in fixed locations outside the template, corresponding to different captions. This categorization focuses on the causes of memorization: MV results from overfitting the one-to-one mapping between image and caption in the training set, while many-to-many mappings lead to RV and TV. However, the concepts of Local Memorization (LM) and Global Memorization (GM) differ, focusing on whether the generated image exactly replicates an entire training image or just a portion of it, regardless of the causes of memorization. RV and TV can be seen as subsets of LM, as there are additional causes of LM beyond the “many-to-many mapping overfitting” that remain unidentified. For instance, in Fig. 8, the image in the third column is classified as MV by Webster (2023) due to the one-to-one mapping of the training image-caption, rather than many-to-many. However, it only memorizes a local portion of the image in the last column. This suggests that a pre-trained model may occasionally exhibit creativity when recalling overfitted memories, which can also lead to local memorization.
@@
Carlini et al. (2023) detects memorization by calculating the Euclidean distance between generated images and training images, recommending re-training on de-duplicated data as a mitigation strategy. Daras et al. (2023) suggests mitigating memorization by training models on corrupted data. Somepalli et al. (2023b) proposes to mitigate memorization by injecting randomness into text prompts, such as adding random tokens or noise, to break the link between memorized prompts and their corresponding images. Chen et al. (2024) employs a nearest neighbor search, similar to Carlini et al. (2023), to identify potential memorized outputs and then selectively applies guidance methods to instruct these generations away from the memorized training images, demonstrating outstanding mitigation performance. Recently, Wen et al. (2024) proposed an efficient method for detecting memorization during the inference phase of diffusion models. This approach uses the magnitude of text-conditional predictions for prompt detection, enabling high accuracy from the initial inference step, and triggers a mitigation strategy to optimize the prompt embedding when memorization is detected. Both the proposed detection and mitigation strategies have achieved state-of-the-art results. A most recent work (Ren et al., 2024) utilized cross-attention patterns to design detection and mitigation strategies, achieving performance comparable to Wen et al. (2024). The key insight is that the cross-attention scores for memorized prompts are less concentrated across all tokens in a prompt during multiple inference steps. This lack of concentration is quantified using entropy, which captures the distribution of cross-attention at the prompt-level. While both leverage cross-attention, it is important to note that our bright ending cross-attention (BE) differs from the cross-attention entropy approach regarding objectives, insights, inputs, and the final computed statistic. Specifically, different from Ren et al. (2024)’s objective of leveraging cross-attention for designing detection and mitigation strategies, BE also aims to achieve a new task of precisely and efficiently locating localized memorization regions. Also, BE does not rely on the concentration patterns of all tokens during multiple steps as the insight and inputs but instead uses the raw final step attention score of the end token. Also, BE employs the raw image-patch level distribution rather than the prompt-level distribution as the computed statistic, serving as a local memorization mask.
7 Conclusion
@@
In this paper, we identify a performance gap in existing tasks of measuring, detecting, and mitigating memorization in diffusion models, particularly in cases where the generated images memorize only parts of the training images. To address this, we introduce a new localized view of memorization in diffusion models, distinguishing between global and local memorization, with the latter being a more meaningful and generalized concept. To facilitate investigation from a local perspective, we propose a new task: extracting localized memorized regions. We propose leveraging the newly observed ‘bright ending’ (BE) phenomenon to precisely and efficiently accomplish this task, resulting in a local memorization mask. Through extensive experiments, we demonstrate that the mask generated by this new task can also be used to enhance performance in existing tasks. Specifically, we propose a simple yet effective strategy to incorporate the BE local memorization mask into existing frameworks, refining them to adopt a local perspective. This approach reduces the performance gap associated with local memorization and achieves new state-of-the-art results. The improved results not only highlight our improvements in existing tasks but also serve as extrinsic validation of BE’s effectiveness in the new task of local memorization mask extraction, underscoring the significant contribution of BE.
8 Acknowledgement
This work was supported in part by the Australian Research Council under Projects DP210101859 and FT230100549.